OpenStudy (anonymous):

A library uses 5 character code to identify the different types of books. The 1st character can be X or Y. The 2nd, 3rd, and 4th can all be 0 through 9. The 5th is A, B , C, D. Examples are : X004C, Y155A, and Y987D. How many different types of books can be identified using this code?

OpenStudy (anonymous):

counting principle yet again two choices for first character, 10 for second 10 for third, 10 for 4th 4 for 5th multiply all together and get \[2\times 10\times 10\times 10\times 4\]